OpenStudy (anonymous):

Algebra 1 help? It has to deal w/ Box and Whisker plots which I have a hard time understanding. :/ A medal and a fan will be given for the help!

OpenStudy (anonymous):

@poopsiedoodle may I have some help?

OpenStudy (anonymous):

Oh 0___0 ok. Hold on:

OpenStudy (ikram002p):

wats ur qs :D

OpenStudy (anonymous):

The third column is Class C.

OpenStudy (poopsiedoodle):

o.o I guess the second row is B and the third is C?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

@ikram002p what do you think? :-)

OpenStudy (ikram002p):

ya the third is C ok ill help u with this :)

OpenStudy (anonymous):

Thank you! <3 @ikram002p

OpenStudy (anonymous):

Thank your for the help! :)

OpenStudy (ikram002p):

ohkk but that was helpfull :D

OpenStudy (anonymous):

Please don't go @ikram002p ! I still don't understand!

OpenStudy (ikram002p):

no no lol i wont go XD

OpenStudy (anonymous):

Oh thank goodness, I need all the help I can get! XD

OpenStudy (ikram002p):

ok the mean is the sum of valus devided bye there count did u got this ?

OpenStudy (anonymous):

I got for Class A the mean is 79, correct?

OpenStudy (anonymous):

And now I got what poopsie got; 85 and 1/3 for the mean.

OpenStudy (ikram002p):

i got 80.86 1213/15

OpenStudy (anonymous):

Hmm...how did you get that?

OpenStudy (ikram002p):

100+90+94+100+75+30+94+72+85+87+42+100+90+94+60 =1213

OpenStudy (anonymous):

Oh! It's Class A and B i'm supposed to find the mean and Standard Deviation! :D

OpenStudy (anonymous):

That's why our second answer is different!

OpenStudy (ikram002p):

hhh i knw lol but it will help u to get the idea XD

OpenStudy (anonymous):

Got it! :)

OpenStudy (ikram002p):

so for A wat u got ?

OpenStudy (anonymous):

For A I got the mean for Class A is 79. Alright, now how do I get the Standard Deviation?

OpenStudy (ikram002p):

for standerd deviation u need to find how much the value is away from the mean ill show u how ok ?

OpenStudy (anonymous):

Ok!

OpenStudy (ikram002p):

first find X'-X for all value , s.t X' is the mean

OpenStudy (anonymous):

Ok so 79 - ? where would I find the X?

OpenStudy (ikram002p):

well work on colum A 80+85+87+92+86+90+94+100+75+85+38+72+94+62=1140 X'=1140/15=76

OpenStudy (anonymous):

You forgot the 45 XD

OpenStudy (anonymous):

Ok so I have 79 and...

OpenStudy (ikram002p):



OpenStudy (anonymous):

Oh!!! WAIT...I SUBTRACT THESE NUMBERS FROM THE MEAN RIGHT?!!?